OpenStudy (anonymous):

What is the 12th term of the arithmetic sequence where a1 = 5 and a6 = 20 ? 29 42 38 34

OpenStudy (anonymous):

d = (20-5)/5 = 3 a1 + 11(3) = 5 + 33 = 38